# Fractal Graph Contrastive Learning

### v2

@@ 3.5 Practical Implementation
$\mu_{n}\sim\mathcal{N}\bigl{(}0,\tau^{2}(D_{n})\bigr{)},\quad\nu_{nk}\sim\mathcal{N}\bigl{(}|\dim(G_{n})-\dim(G_{k})|,\,\tau^{2}(D_{n})+\tau^{2}(D_{k})\bigr{)},$
where $D_{n}=\operatorname{diam}(G_{n})$ and $\hat{\sigma}\approx 0.1$ is the pilot‐estimated residual scale. The fractal loss then reads $\ell_{n}^{\text{fractal}}=-\log\frac{\exp\bigl{(}\mathrm{sim}(\mathbf{z}_{n},\mathbf{z}_{n}^{(\mathscr{R})})/\tau+\alpha\mu_{n}\bigr{)}}{\displaystyle\sum_{k\neq n}\exp\bigl{(}\mathrm{sim}(\mathbf{z}_{n},\mathbf{z}_{k}^{(\mathscr{R})})/\tau+\alpha\nu_{nk}\bigr{)}}.$
Implementation details. During training we first compute (or cache) each graph diameter $D_{i}$, then form the similarity matrix $\mathbf{S}=[\mathrm{sim}(\mathbf{z}_{i},\mathbf{z}_{j}^{(\mathscr{R})})]$ and augment it with a Gaussian matrix whose entrywise statistics obey the diameter–controlled variance above:
$\mathbf{S}^{\ast}=\mathbf{S}+\alpha\mathbf{G},\qquad\mathbf{G}_{ij}\sim\begin{cases}\mathcal{N}\bigl{(}0,\tau^{2}(D_{i})\bigr{)},&i=j,\\[4.0pt] \mathcal{N}\bigl{(}|\dim(G_{i})-\dim(G_{j})|\ ,\ \tau^{2}(D_{i})+\tau^{2}(D_{j})\bigr{)},&i\neq j.\end{cases}$
@@ 4.1 Setup
We validate FractalGCL on unsupervised representation learning tasks using six widely-adopted datasets from TUDataset morris2020tudataset : NCI1, MUTAG, PROTEINS, D&D, REDDIT-BINARY(REDDIT-B), and REDDIT-MULTI-5K(REDDIT-M5K). We adopt a 2-layer GIN as the encoder, and a sum pooling is used as the readout function; renormalisation adopts greedy box-covering with radius 1, dimension weight $\alpha=0.1$, and temperature $\tau=0.4$. Models are first trained with Adam on the unlabeled data only. After that, a non-linear SVM classifier is used to evaluate the graph representations. Accuracy is reported under 10-fold cross-validation. The experiments are repeated 5 times to report the mean and standard deviation. We conduct our experiments on an Ubuntu machine with one 40GB NVIDIA A100 GPU.
@@ 4.2 Main Results
GAE kipf2016variational 74.36 $\pm$ 0.24 72.87 $\pm$ 6.84 70.51 $\pm$ 0.17 74.54 $\pm$ 0.68 87.69 $\pm$ 0.40 33.58 $\pm$ 0.13 68.93
graph2vec narayanan2017graph2vec 73.22 $\pm$ 1.81 83.15 $\pm$ 9.25 73.30 $\pm$ 2.05 70.32 $\pm$ 2.32 75.48 $\pm$ 1.03 47.86 $\pm$ 0.26 70.56
DGI velickovic2019deep 74.86 $\pm$ 0.26 66.49 $\pm$ 2.28 72.27 $\pm$ 0.40 75.78 $\pm$ 0.34 88.66 $\pm$ 0.95 53.61 $\pm$ 0.31 71.95
InfoGraph sun2019infograph 76.20 $\pm$ 1.06 89.01 $\pm$ 1.13 74.44 $\pm$ 0.31 72.85 $\pm$ 1.78 82.50 $\pm$ 1.42 53.46 $\pm$ 1.03 74.74
GraphCL graphcl 77.87 $\pm$ 0.41 86.80 $\pm$ 1.34 74.39 $\pm$ 0.45 78.62 $\pm$ 0.40 89.53 $\pm$ 0.84 55.99 $\pm$ 0.28 77.20
ContextPred hu2019strategies 73.00 $\pm$ 0.30 71.75 $\pm$ 7.34 70.23 $\pm$ 0.63 74.66 $\pm$ 0.51 84.76 $\pm$ 0.52 51.23 $\pm$ 0.84 70.94
JOAO you2021graph 78.07 $\pm$ 0.47 87.35 $\pm$ 1.02 74.55 $\pm$ 0.41 77.32 $\pm$ 0.54 85.29 $\pm$ 1.35 55.74 $\pm$ 0.63 76.39
JOAOv2 you2021graph 78.36 $\pm$ 0.53 87.67 $\pm$ 0.79 74.07 $\pm$ 1.10 77.40 $\pm$ 1.15 86.42 $\pm$ 1.45 56.03 $\pm$ 0.27 76.66
SimGRACE xia2022simgrace 79.12 $\pm$ 0.44 89.01 $\pm$ 1.31 74.03 $\pm$ 0.09 77.44 $\pm$ 1.11 89.51 $\pm$ 0.89 55.91 $\pm$ 0.34 77.50
RGCL li2022let 78.14 $\pm$ 1.08 87.66 $\pm$ 1.01 75.03 $\pm$ 0.43 78.86 $\pm$ 0.48 90.34 $\pm$ 0.58 56.38 $\pm$ 0.40 77.74
DRGCL ji2024rethinking 78.70 $\pm$ 0.40 89.50 $\pm$ 0.60 75.20 $\pm$ 0.60 78.40 $\pm$ 0.70 90.80 $\pm$ 0.30 56.30 $\pm$ 0.20 78.15
GradGCL li2024gradgcl 79.72 $\pm$ 0.53 88.46 $\pm$ 0.98 74.89 $\pm$ 0.39 78.95 $\pm$ 0.47 90.45 $\pm$ 1.06 56.20 $\pm$ 0.31 78.11
FractalGCL 80.50 $\pm$ 0.16 91.71 $\pm$ 0.23 75.85 $\pm$ 0.40 80.14 $\pm$ 0.12 88.13 $\pm$ 0.13 56.45 $\pm$ 0.14 78.80
Table 1 reports the accuracy of the downstream graph classification task on six benchmark datasets. FractalGCL achieves the highest average score (78.80%), outperforming the strongest baseline (GradGCL, $78.11\%$) by 0.69 pp. It ranks first on five of the six datasets—NCI1, MUTAG, PROTEINS, D&D, and REDDIT-MULTI-5K. The most strongly fractal benchmark D&D exhibits a big margin (+1.19 pp), which is consistent with our hypothesis that fractal-aware augmentations and loss provide greater benefit when the underlying graphs display pronounced self-similarity. These results confirm that injecting the fractal structure into a graph contrastive learning not only matches but often exceeds the performance of carefully tuned augmentation-based methods, while retaining the same encoder capacity and training budget. It is worth noting that FractalGCL performs noticeably worse on the Reddit-Binary dataset. This dataset contains the lowest proportion of strongly fractal graphs among all six benchmarks (see Appendix A), which aligns with the theoretical expectation that our method excels when fractal structure is prevalent.
@@ 4.3 Evaluation on Practical Scenarios
To assess the real-world applicability of FractalGCL, we followed network collection way in zhai2025heterogeneous to construct urban road graphs for Chicago, New York and San Francisco. We then randomly sampled square sub-graphs from each complete road network. The downstream task predicts the traffic-accident severity of each area, following zhao2024exploring . Full experimental details and results are provided in Appendix B.
@@
Task City DGI InfoGraph GCL JOAO SimGRACE FractalGCL
total_accidents_high Chicago 54.91$\pm$10.75 56.54$\pm$11.85 63.12$\pm$13.36 55.86$\pm$12.21 62.75$\pm$13.50 64.60$\pm$13.32
SF 76.45$\pm$14.47 78.74$\pm$13.60 80.06$\pm$13.32 79.75$\pm$13.61 80.40$\pm$13.47 80.89$\pm$12.92
NY 51.51$\pm$7.62 51.85$\pm$8.82 55.83$\pm$12.38 51.10$\pm$11.01 52.27$\pm$12.06 68.39$\pm$13.84
accident_volume_level Chicago 43.09$\pm$11.96 42.77$\pm$12.15 46.90$\pm$13.71 43.15$\pm$12.55 46.26$\pm$13.34 48.83$\pm$13.68
SF 55.85$\pm$11.48 58.07$\pm$10.67 58.31$\pm$10.86 58.40$\pm$10.61 58.43$\pm$11.10 58.10$\pm$11.37
NY 37.14$\pm$10.59 35.72$\pm$10.21 39.22$\pm$13.34 35.86$\pm$12.14 36.01$\pm$13.09 50.17$\pm$14.79
risk_level Chicago 34.19$\pm$7.29 34.02$\pm$7.68 40.49$\pm$12.05 33.90$\pm$9.30 37.74$\pm$11.12 43.95$\pm$12.63
SF 39.47$\pm$11.15 40.98$\pm$10.84 41.26$\pm$11.59 41.76$\pm$12.48 41.41$\pm$11.76 42.34$\pm$12.20
NY 41.46$\pm$11.21 42.06$\pm$11.21 47.30$\pm$13.63 41.40$\pm$12.38 42.80$\pm$13.05 57.24$\pm$14.23
Average 48.23 48.97 52.50 49.02 50.90 57.17
Table 2 summarizes three downstream tasks across three cities, yielding nine classification settings in total. FractalGCL attains the highest accuracy in eight of the nine settings and lifts the overall average to 57.17%, an impressive 6.27% lead over the next-best model, SimGRACE (50.90%). We attribute this gain to the strongly fractal nature of urban networks, which makes a fractal-aware approach especially effective; further details appear in Appendix B and publicly available code.
@@ 4.6 Parameter Analysis
Figure 6 shows that FractalGCL is robust to reasonable hyper-parameter changes: accuracy varies within about one percentage point across all tested settings. A moderate fractality filter ($R^{2}\approx 0.9$) and a small dimension weight ($\alpha\leq 0.3$) already captures most of the gain, while larger penalties or very loose filters begin to erode performance. Batch size has little impact, confirming that the method scales smoothly without delicate tuning of training throughput.
5 Related Works
Fractal Geometry for Graphs. Fractal geometry interfaces with graph theory most tangibly through the study of complex networks watts1998collective ; barabasi1999emergence ; song2005self . Analytical models of genuinely fractal graphs remain comparatively scarce. Recently, the Iterated Graph Systems framework has gained attention for its rigorous yet flexible recursive construction of fractal graphs li2024scale ; neroli2024fractal . However, applications of fractal geometry within graph representation learning are still rare.
Graph Contrastive Learning. Graph contrastive learning comprises several crucial stages, among which graph data augmentation assumes a pivotal role, yet it is rendered particularly challenging by the intricate non-Euclidean characteristics inherent in graph topologies gclsurvey1 . Existing graph data augmentation techniques dgi ; graphcl ; joao ; gcc ; rgcl ; gcs ; graphdiff1 ; ji2024rethinking have achieved notable progress. However, they often fall short in adequately preserving the structural similarity between positive pairs, which arises from the inherent difficulty in precisely leveraging complex topological features.
@@ 6 Conclusion
A limitation of FractalGCL is that, on the graphs that do not exhibit fractal structure, its empirical performance may lag behind the results obtained on datasets with pronounced fractal patterns. However, we believe that FractalGCL is not only a compelling demonstration of how fractal geometry can be integrated into machine learning, but also a foundational cornerstone for future research on fractal networks. We intend to continue investigating this line of research in future studies.
@@ Appendix A Proofs
A rigorous mathematical approach to define and analyse “fractal graphs” relies on viewing a graph as a metric space and studying its scaling limit in the sense of Gromov-Hausdorff topology. Concretely, in this section, let $G=(V,E)$ be a simple, connected infinite graph with its shortest-path metric. A sequence of such graphs $(\widehat{G}^{n})$ is said to converge to a limiting metric space $\widehat{G}^{\infty}$ if
$\lim_{n\to\infty}\mathrm{d}_{\mathrm{GH}}\bigl{(}\widehat{G}^{n},\widehat{G}^{\infty}\bigr{)}=0,$
where $\mathrm{d}_{\mathrm{GH}}$ is the Gromov–Hausdorff distance. If the limit $\widehat{G}^{\infty}$ exhibits fractal behaviour, then the original sequence $(G^{n})$ is often viewed to possess fractality in a limiting sense.
Although this framework is theoretically well-founded and widely studied in the context of metric geometry and fractal analysis, it typically introduces extensive technical details. In real-world problems involving large-scale networks (e.g., deep neural architectures, biological networks, or social graphs), a full treatment of Gromov–Hausdorff convergence can be unnecessarily complex. Consequently, the present work uses the notion of “infinite graphs” and “fractal-like structures” primarily as an intuitive and useful abstraction of multi-scale patterns, rather than relying on a strict Gromov-Hausdorff scaling limit argument. Readers interested in the detailed mathematical background are referred to [4, 5, 26] for further discussion.
@@
$\dim_{B}(G):=\lim_{L/\mathrm{diam}(G)\to 0}\frac{\log N_{L}(G)}{-\,\log\,\frac{L}{\mathrm{diam}(G)}},$
@@
If $\dim_{B}(G)$ is both finite and strictly positive, we say that $G$ possesses a fractal property, and we refer to $G$ as a Minkowski graph.
@@
Let $G$ be a finite graph. Denote by $d$ the diameter of $G$, i.e., the maximum shortest-path length between any two nodes. If $d\leq 9$, the scale is considered too small for fractal analysis and the algorithm terminates without output.
For a scale parameter $l\in\{1,2,\ldots,L_{\max}\},\quad\text{with }L_{\max}=\lfloor d/2\rfloor,$ define the covering radius as $r=\lfloor l/2\rfloor.$ At each scale $l$, a greedy box-covering algorithm is applied to compute the number of boxes $N_{B}(l)$ needed to cover $G$:
  • Even $l$: Let $U$ be the set of uncovered nodes. Repeatedly, select a node $v\in U$ whose ball $B(v,r)$ (i.e., the set of nodes within distance $r$ from $v$) covers the maximum number of nodes in $U$. Remove $B(v,r)$ from $U$ and count one box. Continue until $U$ is empty.
  • Odd $l$: First, attempt to select a pair of adjacent nodes $(v,w)$ from $U$ as dual centres. Define the covering set as $B(v,r)\cup B(w,r)$. If such a pair exists that maximizes the coverage of $U$, remove the union from $U$ and count one box. If no suitable pair is found, revert to the single-centre method as in the even case.
Thus, for each $l$, one obtains a data point $(l,N_{B}(l))$. Assuming that the box-covering number obeys a power-law relation $N_{B}(l)\sim l^{-\dim_{B}(G)},$ where $\dim_{B}(G)$ is the box (fractal) dimension, taking logarithms yields $\log N_{B}(l)=-\dim_{B}(G)\log l+b.$ By performing a least-squares linear regression on the data points $\{(\log l,\,\log N_{B}(l))\},$ we fit the model $y=mx+b,$ and define the fractal dimension as
$\dim_{B}(G)=-m.$
The coefficient of determination, $R^{2}$, quantifies the goodness-of-fit of the log-log linear model; values of $R^{2}$ close to 1 indicate a strong power-law (self-similar) behaviour of the graph.
The algorithm outputs the pair $(R^{2},\,\dim_{B}(G))$ as the evaluation result of the graph’s fractal properties.
Algorithm 2 Three–Step Random-Centre Renormalisation
Input. Graph $G$ with adjacency $A$, radius $r$. Output. Renormalised graph $\mathscr{R}(G)$.
  1. 1.
    Random centre selection. While vertices remain, uniformly pick a centre $u$, and form a super node $U=\{\,v\mid(\sum_{i=1}^{r}A^{\,i})_{uv}>0\}\cup\{u\}$; delete $U$ from the pool.
  2. 2.
    Assignment matrix. Record all super nodes in a binary matrix $S$ with $s_{ij}=1$ iff vertex $j$ belongs to super node $i$.
  3. 3.
    Graph reconstruction. Compute $A_{\text{super}}=SAS^{\top}$ and define $\mathscr{R}(G)$ on this adjacency.
@@
$\dim_{B}(G)$ $=\lim_{L/\mathrm{diam}(G)\to 0}\frac{\log\big{(}cN_{L}(G)\big{)}}{-\,\log\,\frac{L}{\mathrm{diam}(G)}}$
@@
If $\Delta_{1}<\Delta_{2}$ and $s_{1}-s_{2}\leq\varepsilon$, for $\varepsilon:=\tau\,\alpha\bigl{(}\Delta_{2}-\Delta_{1}\bigr{)}$, then the fractal-weighted InfoNCE losses satisfy

### v3

@@ 3.5 Practical Implementation
$\mu_{n}\sim\mathcal{N}\bigl(0,\kappa^{2}(D_{n})\bigr),\quad\nu_{nk}\sim\mathcal{N}\bigl(|\dim_{\mathrm{B}}(G_{n})-\dim_{\mathrm{B}}(G_{k})|,\kappa^{2}(\mathrm{diam}(G_{n}))+\kappa^{2}(\mathrm{diam}(G_{k}))\bigr),$
where $\hat{\sigma}\approx 0.1$ is the pilot‐estimated residual scale. The fractal loss then reads $\ell_{n}^{\text{fractal}}=-\log\frac{\exp\bigl(\mathrm{sim}(\mathbf{z}_{n},\mathbf{z}_{n}^{(\mathscr{R})})/\kappa+\alpha\mu_{n}\bigr)}{\displaystyle\sum_{k\neq n}\exp\bigl(\mathrm{sim}(\mathbf{z}_{n},\mathbf{z}_{k}^{(\mathscr{R})})/\kappa+\alpha\nu_{nk}\bigr)}.$
Implementation details. During training we first compute (or cache) each graph diameter $\mathrm{diam}(G_{i})$, then form the similarity matrix $\mathbf{S}=[\mathrm{sim}(\mathbf{z}_{i},\mathbf{z}_{j}^{(\mathscr{R})})]$ and augment it with a Gaussian matrix whose entrywise statistics obey the diameter–controlled variance above:
$\mathbf{S}^{\ast}=\mathbf{S}+\alpha\mathbf{G},\qquad\mathbf{G}_{ij}\sim\begin{cases}\mathcal{N}\bigl(0,\kappa^{2}(\mathrm{diam}(G_{i})_{i})\bigr),&i=j,\\[4.0pt] \mathcal{N}\bigl(|\dim_{\mathrm{B}}(G_{i})-\dim_{\mathrm{B}}(G_{j})|\ ,\ \kappa^{2}(\mathrm{diam}(G_{i}))+\kappa^{2}(\mathrm{diam}(G_{j}))\bigr),&i\neq j.\end{cases}$
@@
Safe fallback under weak fractality or small diameters. We employ a two-stage gate: if the graph diameter $\mathrm{diam}(G)\leq 9$ (for which box-dimension estimation is not meaningful) or the fractality is insufficient $R^{2}<\theta$ with default $\theta=0.9$, we disable the renormalised view and the fractal weighting by setting $\alpha=0$, and retain only standard GCL local augmentations (e.g., node dropping). In this case the positive-pair weight is $\exp(\alpha\Delta(G))=1$, so the objective and its gradients reduce exactly to InfoNCE, and the method strictly degenerates to the GCL baseline in these regimes. Consequently, regardless of dataset fractality, the worst-case performance is at least as good as the corresponding GCL baseline. See Section 4.7 and Appendix C for parameter analysis.
@@ 4.1 Setup
We validate FractalGCL on unsupervised representation learning tasks using six widely-adopted datasets from TUDataset (Morris et al., 2020): NCI1, MUTAG, PROTEINS, D&D, REDDIT-BINARY(REDDIT-B), and REDDIT-MULTI-5K(REDDIT-M5K). We adopt a 2-layer GIN as the encoder, and a sum pooling is used as the readout function; renormalisation adopts greedy box-covering with radius 1, dimension weight $\alpha=0.1$, and temperature $\tau=0.4$. Models are first trained with Adam on the unlabeled data only. After that, a non-linear SVM classifier is used to evaluate the graph representations. Accuracy is reported under 10-fold cross-validation. The experiments are repeated 5 times to report the mean and standard deviation. We conduct our experiments on an Ubuntu machine with one 40GB NVIDIA A100 GPU.
@@ 4.2 Main Results
GAE (Kipf & Welling, 2016) 74.36 $\pm$ 0.24 72.87 $\pm$ 6.84 70.51 $\pm$ 0.17 74.54 $\pm$ 0.68 87.69 $\pm$ 0.40 33.58 $\pm$ 0.13 68.93 $\pm$ 1.41
graph2vec (Narayanan et al., 2017) 73.22 $\pm$ 1.81 83.15 $\pm$ 9.25 73.30 $\pm$ 2.05 70.32 $\pm$ 2.32 75.48 $\pm$ 1.03 47.86 $\pm$ 0.26 70.56 $\pm$ 2.79
DGI (Velickovic et al., 2019) 74.86 $\pm$ 0.26 66.49 $\pm$ 2.28 72.27 $\pm$ 0.40 75.78 $\pm$ 0.34 88.66 $\pm$ 0.95 53.61 $\pm$ 0.31 71.95 $\pm$ 0.76
InfoGraph (Sun et al., 2019) 76.20 $\pm$ 1.06 89.01 $\pm$ 1.13 74.44 $\pm$ 0.31 72.85 $\pm$ 1.78 82.50 $\pm$ 1.42 53.46 $\pm$ 1.03 74.74 $\pm$ 1.12
GraphCL (You et al., 2020) 77.87 $\pm$ 0.41 86.80 $\pm$ 1.34 74.39 $\pm$ 0.45 78.62 $\pm$ 0.40 89.53 $\pm$ 0.84 55.99 $\pm$ 0.28 77.20 $\pm$ 0.62
ContextPred (Hu et al., 2020) 73.00 $\pm$ 0.30 71.75 $\pm$ 7.34 70.23 $\pm$ 0.63 74.66 $\pm$ 0.51 84.76 $\pm$ 0.52 51.23 $\pm$ 0.84 70.94 $\pm$ 1.69
JOAO (You et al., 2021) 78.07 $\pm$ 0.47 87.35 $\pm$ 1.02 74.55 $\pm$ 0.41 77.32 $\pm$ 0.54 85.29 $\pm$ 1.35 55.74 $\pm$ 0.63 76.39 $\pm$ 0.74
JOAOv2 (You et al., 2021) 78.36 $\pm$ 0.53 87.67 $\pm$ 0.79 74.07 $\pm$ 1.10 77.40 $\pm$ 1.15 86.42 $\pm$ 1.45 56.03 $\pm$ 0.27 76.66 $\pm$ 0.88
SimGRACE (Xia et al., 2022a) 79.12 $\pm$ 0.44 89.01 $\pm$ 1.31 74.03 $\pm$ 0.09 77.44 $\pm$ 1.11 89.51 $\pm$ 0.89 55.91 $\pm$ 0.34 77.50 $\pm$ 0.70
RGCL (Li et al., 2022) 78.14 $\pm$ 1.08 87.66 $\pm$ 1.01 75.03 $\pm$ 0.43 78.86 $\pm$ 0.48 90.34 $\pm$ 0.58 56.38 $\pm$ 0.40 77.74 $\pm$ 0.66
DRGCL (Ji et al., 2024) 78.70 $\pm$ 0.40 89.50 $\pm$ 0.60 75.20 $\pm$ 0.60 78.40 $\pm$ 0.70 90.80 $\pm$ 0.30 56.30 $\pm$ 0.20 78.15 $\pm$ 0.47
GradGCL (Li et al., 2024) 79.72 $\pm$ 0.53 88.46 $\pm$ 0.98 74.89 $\pm$ 0.39 78.95 $\pm$ 0.47 90.45 $\pm$ 1.06 56.20 $\pm$ 0.31 78.11 $\pm$ 0.62
FractalGCL 80.50 $\pm$ 0.16 91.71 $\pm$ 0.23 75.85 $\pm$ 0.40 81.71 $\pm$ 0.57 90.41 $\pm$ 0.72 57.29 $\pm$ 0.59 79.58 $\pm$ 0.45
Table 1 reports the accuracy of the downstream graph classification task on six benchmark datasets. FractalGCL achieves the highest average score (79.58%), outperforming the strongest baseline (GradGCL, $78.15\%$) by 1.43 pp. It ranks first on five of the six datasets—NCI1, MUTAG, PROTEINS, D&D, and REDDIT-MULTI-5K. The most strongly fractal benchmark D&D exhibits a big margin (2.76pp), which is consistent with our hypothesis that fractal-aware augmentations and loss provide greater benefit when the underlying graphs display pronounced self-similarity. These results confirm that injecting the fractal structure into a graph contrastive learning not only matches but often exceeds the performance of carefully tuned augmentation-based methods, while retaining the same encoder capacity and training budget.
@@ 4.3 Evaluation on Practical Scenarios
To assess the real-world applicability of FractalGCL, we followed network collection way in (Zhai et al., 2025) to construct urban road graphs for Chicago, New York and San Francisco. We then randomly sampled square sub-graphs from each complete road network. The downstream task predicts the traffic-accident severity of each area, following (Zhao et al., 2024). Full experimental details and results are provided in Appendix B.
@@
Task City DGI InfoGraph GCL JOAO SimGRACE DRGCL GradGCL FractalGCL
[…]
Table 2 summarizes three downstream tasks across three cities, yielding nine classification settings in total. FractalGCL attains the highest accuracy in eight of the nine settings and lifts the overall average to 57.17%, an impressive 3.85% lead over the next-best model, DRGCL (53.32%). We attribute this gain to the strongly fractal nature of urban networks, which makes a fractal-aware approach especially effective; further details appear in Appendix B and publicly available code.
4.4 Validation of Theorem 3.9
Figure 6: Gaussion validation
We aim to test that the change in the Minkowski dimension after one-step renormalisation $\Delta(G)$ is Gaussian in distribution. On the TUDataset D&D benchmark, we find $\mathrm{mean}\Delta(G)=-0.1084$ and $\mathrm{std}=0.1058$ with $n=1178$; The regression $D^{\prime}=\alpha+\beta D$ yields $\alpha=0.0512$, $\beta=0.9711$ and $R^{2}=0.87$, and $\mathrm{corr}(D,D^{\prime}-D)=0.0214$ $(p=0.68)$. Hence, the experimental evidence supports Theorem 3.9.
@@ 4.7 Parameter Analysis
Figure 7 shows that FractalGCL is robust to reasonable hyper-parameter changes: accuracy varies within about one percentage point across all tested settings. A moderate fractality filter ($R^{2}\approx 0.9$) and a small dimension weight ($\alpha\leq 0.3$) already captures most of the gain, while larger penalties or very loose filters begin to erode performance. Batch size has little impact, confirming that the method scales smoothly without delicate tuning of training throughput. Also See Appendix C.
@@ 5 Conclusion
We believe that FractalGCL is not only a compelling demonstration of how fractal geometry can be integrated into machine learning, but also a foundational cornerstone for future research on fractal networks. We intend to continue investigating this line of research in future studies.
@@ Appendix A Proofs
A rigorous mathematical approach to define and analyse “fractal graphs” relies on viewing a graph as a metric space and studying its scaling limit in the sense of Gromov-Hausdorff topology. Concretely, in this section, let $G=(V,E)$ be a simple, connected infinite graph with its shortest-path metric. A sequence of such graphs $({G}^{n})$ is said to converge to a limiting metric space ${G}^{\infty}$ if
$\lim_{n\to\infty}\mathrm{d}_{\mathrm{GH}}\bigl({G}^{n},{G}^{\infty}\bigr)=0,$
where $\mathrm{d}_{\mathrm{GH}}$ is the Gromov–Hausdorff distance. If the limit ${G}^{\infty}$ exhibits fractal behaviour, then the original sequence $(G^{n})$ is often viewed to possess fractality in a limiting sense.
Although this framework is theoretically well-founded and widely studied in the context of metric geometry and fractal analysis, it typically introduces extensive technical details. In real-world problems involving large-scale networks (e.g., deep neural architectures, biological networks, or social graphs), a full treatment of Gromov–Hausdorff convergence can be unnecessarily complex. Consequently, the present work uses the notion of “infinite graphs” and “fractal-like structures” primarily as an intuitive and useful abstraction of multi-scale patterns, rather than relying on a strict Gromov-Hausdorff scaling limit argument. Readers interested in the detailed mathematical background are referred to (Falconer, 2013; Gromov et al., 1999; Neroli, 2024) for further discussion.
@@
$\dim_{\mathrm{B}}(G):=\lim_{L/\mathrm{diam}(G)\to 0}\frac{\log N_{L}(G)}{-\log\frac{L}{\mathrm{diam}(G)}},$
@@
If $\dim_{\mathrm{B}}(G)$ is both finite and strictly positive, we say that $G$ possesses a fractal property, and we refer to $G$ as a Minkowski graph.
@@
1:Graph $\mathcal{G}$ with node set $\mathcal{V}$ and diameter $d$.
2:Fractality metric $R^{2}$ and box dimension $\dim_{\mathrm{B}}(\mathcal{G})$.
3:if $d\leq 9$ then $\triangleright$ too small for fractal analysis
4:  $R^{2}\leftarrow 0$.
5:  $\dim_{\mathrm{B}}(\mathcal{G})\leftarrow 0$.
6:else
7:  $L_{\max}\leftarrow\lfloor d/2\rfloor$.
8:  $\textit{Array}\leftarrow\varnothing$.
9:  for $l\leftarrow 1$ to $L_{\max}$ do
10:   $r\leftarrow\lfloor l/2\rfloor$.
11:   $\mathcal{V}_{\mathrm{remain}}\leftarrow\mathcal{V}$.
12:   $N_{B}(l)\leftarrow 0$.
13:   if $l$ is even then
14:     while $\mathcal{V}_{\mathrm{remain}}\neq\varnothing$ do
15:      $v\leftarrow\arg\max_{v\in\mathcal{V}_{\mathrm{remain}}}|B(v,r)|$. $\triangleright$ $B(v,r)=\{i\in\mathcal{V}_{\mathrm{remain}}\mid d_{\mathcal{G}}(i,v)\leq r\}$
16:      $\mathcal{V}_{\mathrm{remain}}\leftarrow\mathcal{V}_{\mathrm{remain}}\setminus B(v,r)$.
17:      $N_{B}(l)\leftarrow N_{B}(l)+1$.
18:     end while
[…]
20:     if $\exists v,w\in\mathcal{V}_{\mathrm{remain}}\text{ with }d_{\mathcal{G}}(v,w)=1$ then
21:      $(v,w)\leftarrow\arg\max_{v,w\in\mathcal{V}_{\mathrm{remain}}}|B(v,r)\cup B(w,r)|$.
22:      $\mathcal{V}_{\mathrm{remain}}\leftarrow\mathcal{V}_{\mathrm{remain}}\setminus\big(B(v,r)\cup B(w,r)\big)$.
23:      $N_{B}(l)\leftarrow N_{B}(l)+1$.
[…]
25:      $v\leftarrow\arg\max_{v\in\mathcal{V}_{\mathrm{remain}}}|B(v,r)|$.
26:      $\mathcal{V}_{\mathrm{remain}}\leftarrow\mathcal{V}_{\mathrm{remain}}\setminus B(v,r)$.
27:      $N_{B}(l)\leftarrow N_{B}(l)+1$.
28:     end if
29:   end if
30:   $\textit{Array}\leftarrow\textit{Array}\cup\{(\log l,\ \log N_{B}(l))\}$.
31:  end for
32:  Fit $y=mx+b$ to Array by least squares and compute $R^{2}$.
33:  $\dim_{\mathrm{B}}(\mathcal{G})\leftarrow-m$.
34:end if
35:return $R^{2},\ \dim_{\mathrm{B}}(\mathcal{G})$.
Algorithm 2 Algorithm of random-Centre renormalisation
1:Input: Graph $G$ with its Node Set $\mathcal{V}$ and its Adjacency Matrix AA, radius $r$
2:Output: Renormalised Graph $\mathscr{R}(G)$
3:// Initialization
[…]
6:// Random centre selection
7:while $\mathcal{V}_{\text{remain}}\neq\emptyset$ do
8:  $u\leftarrow$ uniformly pick a node $u$ from $\mathcal{V}_{\text{remain}}$
9:  $U\leftarrow\{i|A_{r}[u][i]>0\}\bigcup\{u\}$
10:  $\mathcal{V}_{\text{remain}}\leftarrow\mathcal{V}_{\text{remain}}-U\ ,\ \mathcal{V}_{\text{super}}\leftarrow\mathcal{V}_{\text{super}}\bigcup\{U\}$
11:  set $A_{r}[i][j]$ to 0, for any $i,j\in U$
12:end while
13:// Assignment matrix
[…]
15:// Graph reconstruction
[…]
17:Define renormalised graph $\mathscr{R}(G)$ with $A_{\text{super}}$ as the adjacency matrix
18:return $\mathscr{R}(G)$
@@
$\dim_{\mathrm{B}}(G)$ $=\lim_{L/\mathrm{diam}(G)\to 0}\frac{\log\big(cN_{L}(G)\big)}{-\log\frac{L}{\mathrm{diam}(G)}}$
@@
If $\Delta(H_{1})<\Delta(H_{2})$ and $s(H_{1})-s(H_{2})\leq\tau\alpha\bigl(\Delta(H_{2})-\Delta(H_{1})\bigr)$, then the fractal-weighted InfoNCE losses satisfy